VIOLATION OF MULLIKEN’S RULE FOR RYDBERG STATES

SERHAN N. ALTUNATA, \underline{STEPHEN L. COY}, ROBERT W. FIELD, Dept. of Chemistry and G.R. Harrison Spectroscopy Laboratory, Massachusetts Institute of Technology, Cambridge, Massachusetts 02139.

The results of ab-initio R-matrix calculations display profound deviations from Mulliken’s rule in the continuum electronic structure of CaF. In the 1 − D effective potential for the electronic motion, a shape resonance induces a crossover from short-range Mulliken-type dynamics to long-range dipole-mediated dynamics. This effect is visible in the scattering matrix as well as the photoionization cross-sections and photoelectron anisotropies. The non-Mulliken behavior can be exploited by l-selective detection of Rydberg dynamics and spectra, revealing the l-decomposition of molecular eigenstates.